OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3)

mathslover (mathslover):

first open the brackets wht do u get ?

OpenStudy (anonymous):

u gett -2x+ 6 => 5 - (x+3)

OpenStudy (anonymous):

ugg this is hard

OpenStudy (anonymous):

now collect that similar terms

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

dont get it can u show me step by ste[p

OpenStudy (anonymous):

so you understand -2x +6 >= 5 -(x+3) distribute the negative across the parenthesis, -2x +6 >=5-x-3 collect the similar terms, so that means bring all x to one side, and constants to the other. 6-5 +3 > = -x +2x 4>= x

OpenStudy (anonymous):

wat do u do after u hav 4 >=x

OpenStudy (anonymous):

well, that's the answer, the solution set to said inequality is all numbers greater than or equal to 4